If r 12 =0.7, r 23 =r 31 =0.6, find the partial correlation coefficient between x 1
and x 2 keeping x 3 as constant. Find the multiple correlation coefficient
R 1.23
1
Expert's answer
2020-11-17T06:52:39-0500

r12.3="\frac{r_{12}-r_{23}r_{13}}{\sqrt{1-r^2_{23}}{\sqrt{1-r^2_{13}}}}=\frac{0.7-0.6*0.6}{\sqrt{1-0.6^2}\sqrt{1-0.6^2}}=\frac{0.34}{0.64}=0.53125"